OpenStudy (anonymous):

the half of radon-222 is 3.8 days. How much of a 100 g sample is left after 15.2 days?

OpenStudy (zpupster):

15.2 days = 15.2/3.8 = 4 half days Then leftover = 100 grams * (0.5)^4 = 6.25 grams (50 grams after 3.8 days, 25 grams after 7.6 days, 12.5 grams after 11.4 days, 6.25 grams after 15.2 days).